# EvaSurf: Efficient View-Aware Implicit Textured Surface Reconstruction on Mobile Devices

### v3

@@ Abstract
Figure 1: Examples of reconstruction results of EvaSurf. Our model can reconstruct high-quality appearance and accurate mesh on various devices in real-time for both synthetic and real-world objects.
@@ II-A Neural Rendering for 3D Reconstruction
More recently, neural rendering technique has been widely adopted in the field of 3D object reconstruction. Incorporating the implicit representations, Neural Radiance Field (NeRF) [9] and its followers [8, 10, 12, 13, 50, 51, 52, 53, 54, 55, 56] achieve high-fidelity rendering results. These methods incorporate ray-marching and ray-sampling mechanisms during the process of neural rendering through an encoded 5D continuous field. The neural network estimates volume densities and RGB values along the sampled points on the marching rays. However, the corresponding geometry reconstructed by these methods is not satisfactory and accurate due to the ambiguity of volume rendering. Besides implicit representation, recent 3D Gaussian splatting [27, 26, 28, 29] involves iterative refinement of Gaussians to reconstruct 3D objects from multiview images, allowing for high-quality rendering of novel views. Although Gaussian-based methods can achieve real-time rendering on PCs, real-time rendering on mobile devices remains a challenge due to the requirement of large disk storage and lack of mobile-end neural network optimization.
To improve the quality of geometry, surface-based rendering methods open up another road for 3D reconstruction. These methods optimize the underlying surface directly, or gradually deform the geometry from volumetric representation into a surface representation. The implicit surface representations are encoded in neural networks such as occupancy networks [57, 11, 58, 59] or neural signed distance functions [11, 58, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 32]. Although these methods are able to generate relatively satisfactory geometry and appearance, they often require excessive runtime and thus become impractical for daily real-time applications. Conversely, our method also adopts neural rendering for the appearance but uses explicit mesh to represent the geometry, which is integrated into an efficient two-stage framework and supports real-time rendering.
@@ II-B Real-time Rendering on Mobile Devices
With the rapid development of modern technology, mobile phones, and other mobile devices have become an essential part of people’s daily lives. Real-time rendering on these devices has emerged as an important research topic. Before the introduction of neural radiance fields, several methods [70, 71, 72] employ a neural texture and a neural renderer to render a scene. Though these methods can achieve photo-realistic rendering in real time, they may not be generalized to arbitrary 3D objects due to their requirements of geometry priors. To facilitate the power of NeRF to render scenes on mobile devices, Hedman et al.  [24] propose a solution that bakes NeRF into sparse neural radiance grids (SNeRG). However, the rendering speed of SNeRG fails to achieve real-time performance, hindering the attainment of interactive rendering rates. In contrast, MobileNeRF [25] adopts textured polygons as the scene representation and leverages the polygon rasterization pipeline to accelerate the rendering process. Although MobileNeRF supports real-time rendering with interactive frame rates, the exported mesh may be inaccurate due to its complex interlacing polygons. Re-ReND [30] and NeRF2Mesh [31] both train a NeRF-like model at the initial stage and then bake the appearance into a texture map and extract the corresponding mesh. Nevertheless, these methods [30, 31, 73] are unable to render scenes with view-dependent effects and can not reconstruct accurate mesh for complex structures and non-Lambertian objects. Bakedsdf [32] utilizes a hybrid SDF-volume representation and bakes the representation into meshes for real-time rendering. However, it tends to learn over-smooth surfaces and cannot handle complex structures. In comparison, our method employs a multi-view supervision module to learn the complex structure and prevent the model from over-smoothing. We also utilize a view-aware encoding in feature space for better renderings.
@@ III Methods
Given a set of images =$\mathcal{I}=\{I_{1},...,I_{n}\}$, we aim to optimize a representation to efficiently synthesize novel views and generate accurate mesh at interactive frame rates. To this end, we design a two-stage differentiable framework that requires much less computation and disk storage so as to make it compatible with mobile devices. As shown in Fig. 2, we first utilize a surface-based model $\mathcal{S}$ to reconstruct an accurate geometry (III-A). In the second stage, we rasterize the initial mesh and introduce a view-aware encoding $L$ to embed the view-dependent information into an implicit texture $f$ (III-B). Finally, we jointly optimize the implicit texture $f$ and a neural shader, i.e. , a lightweight MLP, which takes the view-embedded implicit feature $f_{l}$ as input and directly outputs the rendered images $\hat{I}$ (III-C).
@@ III-A Efficient Explicit Geometry Reconstruction
As the base of the reconstruction pipeline, an accurate geometry is crucial for the following high-quality rendering. To overcome the challenge of coarse geometry faced by the traditional NeRFs, our model first trains a surface-based model $\mathcal{S}$ that follows the pipeline of NeuS [61] and leverages a set of signed distance function (SDF) to directly estimate the explicit geometry (mesh).
@@
On one hand, SDF-based models like NeuS [61] tend to generate meshes that have geometric ambiguities, resulting in inaccurate shape. Particularly, concave surfaces are easily interpreted as convex oversmooth surface using such methods. On the other hand, multi-view stereo method can reconstruct accurate concave shapes, but it is prone to generate incomplete meshes. Therefore, to improve the quality of the geometry, we introduce a multi-view supervision (MVS) module to supervise the SDF-based geometry learning. The MVS module first reconstructs the shape using OpenMVS [74] and renders depth images. We then supervise the surface-based model using the rendered depth images. The MVS module can prevent our model from generating over-smooth surfaces and encourage it to learn complex structures. Overall, for an explicit geometry $G$ represented by triangle mesh with vertex positions $V$, the objective is:
@@
$\mathcal{L}_{s}=w_{rgb}L_{rgb}+w_{eikonal}L_{eikonal}+w_{mask}L_{mask}$ (3)
The weight for each loss is determined by default as $w_{rgb}=1.0$, $w_{eikonal}=0.1$, $w_{mask}=0.1$ in experiments.
On one hand, using SDF-based models like Neus [61] can reconstruct a solid mesh with details. However, such methods tend to generate a mesh that lacks smoothness and suffers from geometric ambiguities, resulting in inaccurate shape. On the other hand, using multi-view stereo reconstruction method can reconstruct a smoother mesh with an accurate representation of convex and concave shapes. Nevertheless, such methods are prone to generate a damaged and incomplete mesh. Therefore, to enable the surface-based model to learn the accurate shape of complex structures, we further supervise the model by employing a multi-view supervision module (MVS). Specifically, in the multi-view supervision module, we first employ the open multi-view stereo reconstruction library (OpenMVS) to reconstruct the shape and render depth images. We then supervise the surface-based model using the rendered depth images and the normals calculated from the depths. By employing this additional module, we introduce the depth priors to the surface-based model, encouraging the model to capture accurate shape information while reconstructing a solid and smooth surface. The effectiveness of this module is demonstrated in the experiments section of the paper. The formulation of $\mathcal{L}_{mvs}$ is demonstrated as follows:
$\mathcal{L}_{mvs}=w_{depth}L_{depth}+w_{normal}L_{normal}$ (4)
The weight is determined by default as $w_{depth}=0.3$, $w_{normal}=0.04$ in experiments.
@@ III-B View-aware Implicit Texture
With the introduction of the view-aware encoding, our method introduces the view-dependent information in the feature space instead of the 3D point space. Specifically, the view-dependent information is represented as the concentration of the values in the texture map. By assigning different weights to each dimension of the texture, we model view dependence with different fidelity and thus encourage the texture to learn the view-dependent information in the scene. For example, a cosine value of 1 ($\cos(0)$) corresponds to a frontal view, with the attention of values lying at the front of the implicit texture. In this way, our method is able to generate a range of view-dependent effects and improve the quality of the rendered images.
@@ III-D Differentiable Training
Similar to [75], we adopt the differentiable rendering technique for the training procedure. As depicted in Fig. 2, during the forward-pass procedure, our method renders a 2D image given the explicit geometry and the learnable implicit texture, which altogether affect the shape and appearance of a 3D object. As for the backward-pass stage, the gradient of the loss with respect to the input shape and appearance factors is computed given the gradient of the loss function defined on the output image pixels. Here, we use the L2 loss as the reconstruction loss:
@@
where $(x,y,z,w)$ is the homogeneous coordinates input from the world space and $(u,v)$ is the barycentric coordinates from the image space.
@@ IV-A Experiment Setups
Network Architecture For the network architecture of our surface-based model, we first use TCNNHashGrid from [54] with progressive levels to embed the points in 3d space. In the hierarchy of the grids, we use 16 levels from 16 to 4096, where each levels has 4 channels. The SDF MLP in our model has 2 layers with 64 hidden units. As for the texture network, it is noteworthy that unlike other SDF-based method the texture network in our surface-based model mainly serves to aid the geometry learning instead of producing the final rendering. Learning geometry from multi-view inputs can be a challenging task, so we adopt multi-head texture network during the pipeline. To be more specific, we adopt two texture networks during the geometry learning. We use positional encoding to embed the positions and spherical harmonics encoding to embed the view directions in the first texture network. In the second network, we use positional encoding for both the positions and view-directions. The texture network with multi texture heads accomplishes the task of coarse-to-fine learning. The first texture MLP has 2 layers with 64 hidden units and the second has 3 layers with 64 hidden units.
@@ IV-B Comparisons on Synthetic Data
Geometry Quality. To evaluate the geometry quality, we compare the exported meshes of our method and previous methods [61, 44, 25, 30, 31, 26, 32, 65]. We choose bi-directional Chamfer Distance as the evaluation metric. Nevertheless, due to the fact that the ground truth meshes may not accurately represent surface meshes, we cast rays and sample 2.5 million points from the intersections between the rays and the surfaces in each scene following the evaluation process of [31]. The quantitative results are demonstrated in Table I.Compared with previous implicit methods for surface reconstruction [61, 44, 65], our method achieves superior mesh quality while supporting real-time rendering on mobile devices. Compared with state-of-the-art explicit Gaussian-based methods [26], EvaSurf also reconstructs more accurate geometry for all the scenes while achieve mobile-rendering. Compared with existing methods [25, 30, 31, 32] that allow rendering on mobile devices, our method can generate more accurate mesh for the scenes with view-dependent effects.
@@
Rendering Quality. To evaluate the performance of rendering quality, we compare our results with the state-of-the-art methods that achieve real-time rendering on mobile devices [30, 31, 25]. We also provide the results of the original NeRF model [9] and SNeRG [24] as reference. We choose PSNR, SSIM, and LPIPS as the evaluation metrics. The experiments are conducted on the NeRF synthetic dataset [9]. As shown in Table II, our method generates comparable results with previous methods. It is worth noting that our proposed method yields superior outcomes compared to previous approaches in terms of generating a smooth mesh. Moreover, our approach generates comparable results with methods that prioritize rendering over mesh quality.
@@
For a fair comparison, all the models are trained on a single Tesla-V100 (32G) and the disk storage consists of the uncompressed mesh file (.obj and .mtl), texture file (.png), and the final model file (.json). As shown in Table III, our method consumes smaller disk storage than previous methods. This can be attributed to the utilization of an accurate surface-based model, which prevents the generation of redundant vertices and faces. Additionally, our approach incorporates a lightweight learnable texture in the rendering process, further reducing the overall storage. As for the model training time, our method achieves a much faster training speed than MobileNeRF [25] and Re-ReND [30]. With the incorporation of the learnable implicit texture, our method directly optimizes the final texture required for rendering and bypasses the excessive computation required by these methods [25, 30] to transform the sampling features from the radiance field to the final textures. As for NeRF2Mesh [31], since it utilizes a fast-training NeRF model before obtaining the mesh and texture, it trains slightly faster than our method. However, our model can handle view-dependent surfaces that NeRF2Mesh [31] could not with comparable training time. The rendering speed is measured using iPhone13 and the results are reported in Table. III.
@@ IV-D Ablation Studies
View-aware Encoding. In our real-time 3D object reconstruction method, we optimize a 12-dimensional learnable implicit texture to render high-fidelity images on various devices. To incorporate the view-dependent condition during optimization, we apply view-aware encoding to the original implicit texture. By modeling the view dependence with different fidelity, several view-dependent effects can be rendered by our method. Comparison results on the real-world dataset are displayed in Fig. 8 and the evaluation metrics are also reported in Table V. It could be seen that we obtain better renderings with the set of Gaussians. Since some regions are observable from only a limited range of viewing directions, modeling the view dependence with the same fidelity would lead to an average appearance aggregated from input views. By utilizing the Gaussians in view-aware encoding, we model the view-dependent information with different fidelity. By assigning different weights to each dimension of the texture in the feature space, we enable the model to reflect more realistic results.
@@
TABLE V: Ablation of view-aware encoding. The view-aware encoded model exhibits better view-dependent rendering results.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w $L(f,v_{cos})$ 27.58 0.891 0.208
w/o $L(f,v_{cos})$ 26.69 0.884 0.215
We also provide the detailed evaluation results for the ablation of the view-aware encoding utilized to model view-dependent effects. The PSNR, SSIM, and LPIPS results for each case of the real-world dataset are reported in Table. VI. We also reported the differences $\Delta$PSNR, $\Delta$SSIM and $\Delta$LPIPS in the table to demonstrate the effectivenss. It is noteworthy that by utilizing the view-aware encoding, we obtained better results in all the metrics except for a slight decrease of SSIM and LPIPS for Object_11.
TABLE VI: Per-scene Evaluation for the real-world dataset. We report PSNR $\uparrow$, SSIM $\uparrow$, and LPIPS $\downarrow$ for each case in the real-world dataset. The differences $\Delta$PSNR, $\Delta$SSIM, and $\Delta$LPIPS are also reported.
@@
In our model, we store the 12-dimensional implicit texture as 3 PNG in the rendering package. We then conduct experiments on using 8-dimensional implicit texture and 16-dimensional implicit texture in our model, which are stored as 2 PNG and 4 PNG correspondingly. The evaluation results of PSNR, SSIM, LPIPS, and the final disk storages are reported in Table VII. It could be seen that we obtain improved renderings by increasing the channel number but consume more disk storage. To strike a balance between the rendering quality and disk storage, we choose 12 as the default channel number of the implicit texture. The channel number could be decreased if users have more restrictions on the disk storage and increased if users require better renderings.
TABLE VII: Ablation of texture channel number. We report PSNR, SSIM, LPIPS and the disk storage for reference

### v4

@@ Abstract
Figure 1: Examples of reconstruction results of EvaSurf. Our model can reconstruct high-quality appearance and accurate mesh for both synthetic and real-world objects. EvaSurf also supports real-time rendering on various devices.
@@ II-A Neural Rendering for 3D Reconstruction
More recently, neural rendering technique has been widely adopted in the field of 3D object reconstruction. Incorporating the implicit representations, Neural Radiance Field (NeRF) [9] and its followers [8, 10, 12, 13, 50, 51, 52, 53, 54, 55, 56] achieve high-fidelity rendering results. These methods incorporate ray-marching and ray-sampling mechanisms during the process of neural rendering through an encoded 5D continuous field. The neural network estimates volume densities and RGB values along the sampled points on the marching rays. However, the corresponding geometry reconstructed by these methods is not satisfactory and accurate due to the ambiguity of volume rendering. Besides implicit representation, recent 3D Gaussian splatting [27, 26, 28, 29, 57] involves iterative refinement of Gaussians to reconstruct 3D objects from multiview images, allowing for high-quality rendering of novel views. Although Gaussian-based methods can achieve real-time rendering on PCs, real-time rendering on mobile devices remains a challenge due to the requirement of large disk storage and lack of mobile-end neural network optimization.
To improve the quality of geometry, surface-based methods open up another road for 3D reconstruction. These methods optimize the underlying surface directly, or gradually deform the geometry from volumetric representation into a surface representation. The implicit surface representations are encoded in neural networks such as occupancy networks [58, 11, 59, 60] or neural signed distance functions (SDF) [11, 59, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 32, 71]. These SDF-based models can reconstruct a solid mesh with details. However, such methods tend to generate a mesh that lacks smoothness and suffers from geometric ambiguities, resulting in inaccurate shape. Particularly, concave surfaces are easily interpreted as convex oversmooth surface using such methods.In addition, these methods often require excessive runtime and thus become impractical for daily real-time applications. Multi-view stereo (MVS) reconstruction techniques are then adapted to NeRF [72, 73, 74] for better shape reconstruction. These methods can reconstruct a smoother mesh with an accurate representation of convex and concave shapes. Nevertheless, such methods are prone to generate a damaged and incomplete mesh. Conversely, our method adopts neural rendering for the appearance but uses explicit mesh to represent the geometry. We implement our method into an efficient two-stage framework and supports real-time rendering. Furthermore, we employ an additional supervision module to take advantage of both SDF and MVS methods for better reconstruction.
@@ II-B Real-time Rendering on Mobile Devices
With the rapid development of modern technology, mobile phones, and other mobile devices have become an essential part of people’s daily lives. Real-time rendering on these devices has emerged as an important research topic. Before the introduction of neural radiance fields, several methods [75, 76, 77, 78] employ a neural texture and a neural renderer to render a scene. Though these methods can achieve photo-realistic rendering in real time, they may not be generalized to arbitrary 3D objects due to their requirements of geometry priors.
To facilitate the power of NeRF to render scenes on mobile devices, Hedman et al.  [24] propose a solution that bakes NeRF into sparse neural radiance grids (SNeRG). However, the rendering speed of SNeRG fails to achieve real-time performance, hindering the attainment of interactive rendering rates. In contrast, MobileNeRF [25] adopts textured polygons as the scene representation and leverages the polygon rasterization pipeline to accelerate the rendering process. Although MobileNeRF supports real-time rendering with interactive frame rates, the exported mesh may be inaccurate due to its complex interlacing polygons. Re-ReND [30] and NeRF2Mesh [31] both train a NeRF-like model at the initial stage and then bake the appearance into a texture map and extract the corresponding mesh. Nevertheless, these methods [30, 31, 79] are unable to render scenes with view-dependent effects and can not reconstruct accurate mesh for complex structures and non-Lambertian objects. Bakedsdf [32] utilizes a hybrid SDF-volume representation and bakes the representation into meshes for real-time rendering. However, it tends to learn over-smooth surfaces and cannot handle complex structures. In comparison, our method employs a multi-view supervision module to learn the complex structure and prevent the model from over-smoothing. We also utilize a view-aware encoding in feature space for better renderings.
@@ III Methods
Given a set of images =$\mathcal{I}=\{I_{1},...,I_{n}\}$, we aim to optimize a representation to efficiently generate accurate mesh and support synthesizing novel views at interactive frame rates. To this end, we design a two-stage differentiable framework that requires much less computation and disk storage so as to make it compatible with mobile devices. As shown in Fig. 2, we first utilize a surface-based model $\mathcal{S}$ to reconstruct an accurate geometry (III-A). In the second stage, we rasterize the initial mesh and introduce a view-aware encoding $L$ to embed the view-dependent information into an implicit texture $f$ (III-B). Finally, we jointly optimize the implicit texture $f$ and a neural shader, i.e. , a lightweight MLP, which takes the view-embedded implicit feature $f_{l}$ as input and directly outputs the rendered images $\hat{I}$ (III-C).
@@ III-A Efficient Explicit Geometry Reconstruction
As the base of the reconstruction pipeline, an accurate geometry is crucial for the high-quality rendering. To overcome the challenge of coarse geometry faced by the traditional NeRFs, our model first trains a surface-based model $\mathcal{S}$ that follows the pipeline of NeuS [62] and leverages a set of signed distance functions to directly estimate the explicit geometry.
@@
To enable the surface-based model to learn the accurate shape of complex structures, we introduce a multi-view supervision (MVS) module to supervise the SDF-based geometry learning. The MVS module first reconstructs the shape using the open multi-view stereo reconstruction library OpenMVS [80] and renders depth images. We then supervise the surface-based model using the rendered depth images. The MVS module can prevent our model from generating over-smooth surfaces and encourage it to learn complex structures. Overall, for an explicit geometry $G$ represented by triangle mesh with vertex positions $V$, the objective is:
@@
$\mathcal{L}_{s}=w_{rgb}L_{rgb}+w_{eikonal}L_{eikonal}+w_{mask}L_{mask},$ (3)
where $L_{rgb}$ is the L2 loss, $L_{eikonal}$ is the loss term proposed in NeuS [62] and $L_{mask}$ also follows the definition of NeuS [62]. The weight for each loss is determined by default as $w_{rgb}=1.0$, $w_{eikonal}=0.1$, $w_{mask}=0.1$ in experiments.
In the MVS module, we employ OpenMVS [80] to reconstruct the shape and render depth images. We then supervise the surface-based model using the rendered depth images and the normals calculated from the depths. By employing this additional module, we introduce the depth priors to the surface-based model, encouraging the model to capture accurate shape information while reconstructing a solid and smooth surface. The effectiveness of this module is demonstrated in the experiments section of the paper. The formulation of $\mathcal{L}_{mvs}$ is demonstrated as follows:
$\mathcal{L}_{mvs}=w_{depth}L_{depth}+w_{normal}L_{normal},$ (4)
where $L_{depth}$ is the L1 loss between the depth generated by OpenMVS and SDF-model, and $L_{normal}$ is the L2 loss between the normal derived from the OpenMVS depths and the normal generated by the SDF-mode. The weight is determined by default as $w_{depth}=0.3$, $w_{normal}=0.04$ in experiments.
@@ III-B View-aware Implicit Texture
Specifically, the view-dependent information is represented as the concentration of the values in the texture map. By assigning different weights to each dimension(Chanel) of the texture, we model view dependence with different fidelity and thus encourage the texture to learn the view-dependent information in the scene. Furthermore, our method introduces the view-dependent information in the feature space instead of the 3D point space. Learning the additional view-independent information upon the topological-structured UVs, we captured the anisotropy while preserving the inference efficiency.
@@ III-D Differentiable Training
We adopt the differentiable rendering technique for the training procedure. As depicted in Fig. 2, during the forward-pass procedure, our method renders a 2D image given the explicit geometry and the learnable implicit texture, which altogether affect the shape and appearance of a 3D object. As for the backward-pass stage, the gradient of the loss with respect to the input shape and appearance factors is computed given the gradient of the loss function defined on the output image pixels. Here, we use the L2 loss as the reconstruction loss:
@@
where $(x,y,z,w)$ is the homogeneous coordinates input from the world space and $(u,v)$ is the barycentric coordinates from the image space. Similar to the rendering pipeline of nvdiffrast [81], the mesh is rasterized and the 3D positions are interpolated pixel-wisely in image space. Since we build a topological-structured UVs after rasterization, the appearance from stage 1 can be inherited to stage 2. This approach reduces the required training time for stage 2 by eliminating the need to learn the appearance from scratch.
@@ IV-A Experiment Setups
Network Architecture. For the network architecture of our surface-based model, we first use TCNNHashGrid from [54] with progressive levels to embed the points in 3d space. In the hierarchy of the grids, we use 16 levels from 16 to 4096, where each levels has 4 channels. The SDF MLP in our model has 2 layers with 64 hidden units. As for the texture network, it is noteworthy that unlike other SDF-based method the texture network in our surface-based model mainly serves to aid the geometry learning instead of producing the final rendering. Learning geometry from multi-view inputs can be a challenging task, so we adopt multi-head texture network during the pipeline. To be more specific, we adopt two texture networks during the geometry learning. We use positional encoding to embed the positions and spherical harmonics encoding to embed the view directions in the first texture network. In the second network, we use positional encoding for both the positions and view-directions. The texture network with multi texture heads accomplishes the task of coarse-to-fine learning. The first texture MLP has 2 layers with 64 hidden units and the second has 3 layers with 64 hidden units.
@@ IV-B Comparisons on Synthetic Data
Geometry Quality. To evaluate the geometry quality, we compare the exported meshes of our method and previous methods [62, 44, 25, 30, 31, 26, 32, 66]. We choose bi-directional Chamfer Distance as the evaluation metric. Nevertheless, due to the fact that the ground truth meshes may not accurately represent surface meshes, we cast rays and sample 2.5 million points from the intersections between the rays and the surfaces in each scene following the evaluation process of [31]. The quantitative results are demonstrated in Table I. Compared with previous implicit methods for surface reconstruction [62, 44, 66], our method achieves superior mesh quality while supporting real-time rendering on mobile devices. Compared with state-of-the-art explicit Gaussian-based methods [26, 57], EvaSurf also reconstructs more accurate geometry for all the scenes while achieve mobile-rendering. Compared with existing methods [25, 30, 31, 32] that allow rendering on mobile devices, our method can generate more accurate mesh for the scenes with view-dependent effects.
@@
Rendering Quality. To evaluate the performance of rendering quality, we compare our results with the state-of-the-art methods that achieve real-time rendering on mobile devices [30, 31, 25]. We also provide the results of the original NeRF model [9] and SNeRG [24] as reference. It is noteworthy that we set the texture resolution as 2048 in all models for fair comparison. We choose PSNR, SSIM, and LPIPS as the evaluation metrics. The experiments are conducted on the NeRF synthetic dataset [9]. As shown in Table II, our method generates comparable results with previous methods. It is worth noting that our proposed method yields superior outcomes compared to previous approaches in terms of generating a smooth mesh. Moreover, our approach generates comparable results with methods that prioritize rendering over mesh quality.
@@
For a fair comparison, all the models are trained on a single Tesla-V100 (32G) and the disk storage consists of the uncompressed mesh file (.obj and .mtl), texture file (.png), and the final model file (.json). As shown in Table III, our method consumes smaller disk storage than previous methods. This can be attributed to the utilization of an accurate surface-based model, which prevents the generation of redundant vertices and faces. Additionally, our approach incorporates a lightweight learnable texture in the rendering process, further reducing the overall storage. As for the model training time, our method achieves a much faster training speed than MobileNeRF [25] and Re-ReND [30]. With the incorporation of the learnable implicit texture, our method directly optimizes the final texture required for rendering and bypasses the excessive computation required by these methods [25, 30] to transform the sampling features from the radiance field to the final textures. For gaussian-based methods like 2DGS [26], although they can easily render at interactive framerates, they are not built for end-to-end mobile rendering. To further compare the rendering efficiency, we implement a three.js based renderer for mobile rendering. It can be seen that 2DGS requires excessive disk storages for rendering while our method is more mobile-friendly considering the aspect of storages. As for NeRF2Mesh [31], since it utilizes a fast-training NeRF model before obtaining the mesh and texture, it trains slightly faster than our method. However, our model can handle view-dependent surfaces that NeRF2Mesh [31] could not with comparable training time. The rendering speed is measured using iPhone13 and the results are reported in Table III. It is noteworthy that the reported training time (1-2 h) includes both training time required for stage 1 and stage 2 and the time required for OpenMVS reconstruction.
@@ IV-D Ablation Studies
View-aware Encoding. In our 3D object reconstruction method, we optimize a 12-dimensional learnable implicit texture to render high-fidelity images on various devices. To incorporate the view-dependent condition during optimization, we apply view-aware encoding to the original implicit texture. By modeling the view dependence with different fidelity, several view-dependent effects can be rendered by our method. Comparison results on the real-world dataset are displayed in Fig. 8 and the evaluation metrics are also reported in Table V. It could be seen that we obtain better renderings with the set of Gaussians. Since some regions are observable from only a limited range of viewing directions, modeling the view dependence with the same fidelity would lead to an average appearance aggregated from input views. By utilizing the Gaussians in view-aware encoding, we model the view-dependent information with different fidelity. By assigning different weights to each dimension of the texture in the feature space, we enable the model to reflect more realistic results. Furthermore, we fix the geometry reconstructed from stage 1 and employ the spherical gaussians used in BakedSDF [32] and the MLPs in NeRF2Mesh [31] to further demonstrate the effectiveness of our anisotropy modeling strategies. The results shown in Table V proves that our feature-level Gaussians outperforms other anisotropy modelling strategies.
@@
TABLE V: Ablation of view-aware encoding. The view-aware encoded model exhibits better view-dependent rendering results and is more effective than other anisotropy modeling strategy
. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ w/ $L(f,v_{cos})$ 27.58 0.891 0.208 w/ SG Modeling [32] 27.24 0.889 0.209 w/ MLP Modeling [31] 26.96 0.887 0.212 w/o $L(f,v_{cos})$ 26.69 0.884 0.215
Our method uses feature-level Gaussians to model anisotropy and set the hyperparametr $\alpha$ in Eq. 6 as 0.4 in experiments. This hyperparameter models the variance of the feature-level Gaussians, and the value should be determined carefully to ensure the model can capture appropriate range of anisotropy. Therefore, we present the ablation experiments on $\alpha$ and the rendering quality in Table VI. It can be seen that either too large or too small would lead to inferior rendering results, and we determine the value through experiments and find out that the optimal value lies in around [0.4, 0.45], slightly leaning towards 0.4. Therefore, we chose $\alpha$ to be 0.4 by default.
TABLE VI: Ablation on the hyperparameter $\alpha$. We present the evaluation of PSNR by adjusting the value of $\alpha$
$\alpha$ 0.1 0.2 0.4 0.6 0.8 1.0 1.4
PSNR ($\uparrow$) 26.64 26.85 27.58 27.39 26.42 26.04 25.96
We also provide the detailed evaluation results for the ablation of the view-aware encoding utilized to model view-dependent effects. The PSNR, SSIM, and LPIPS results for each case of the real-world dataset are reported in Table VII. We also reported the differences $\Delta$PSNR, $\Delta$SSIM and $\Delta$LPIPS in the table to demonstrate the effectivenss. It is noteworthy that by utilizing the view-aware encoding, we obtained better results in all the metrics except for a slight decrease of SSIM and LPIPS for Object_11.
TABLE VII: Per-scene Evaluation for the real-world dataset. We report PSNR $\uparrow$, SSIM $\uparrow$, and LPIPS $\downarrow$ for each case in the real-world dataset. The differences $\Delta$PSNR, $\Delta$SSIM, and $\Delta$LPIPS are also reported.
@@
Implicit Texture. In our model, we store the 12-dimensional implicit texture as 3 PNG in the rendering package. We then conduct experiments on using 8-dimensional implicit texture and 16-dimensional implicit texture in our model, which are stored as 2 PNG and 4 PNG correspondingly. The evaluation results of PSNR, SSIM, LPIPS, and the final disk storages are reported in Table VIII. It could be seen that we obtain improved renderings by increasing the channel number but consume more disk storage. To strike a balance between the rendering quality and disk storage, we choose 12 as the default channel number of the implicit texture. The channel number could be decreased if users have more restrictions on the disk storage and increased if users require better renderings.
TABLE VIII: Ablation of texture channel number. We report PSNR, SSIM, LPIPS and the disk storage for reference
@@
In addition, we use 2048 as the texture resolution by default, and we conducted ablation experiment on the texture resolution using 1024, 2048 and 4096 as the resolution. The evaluation results of PSNR, disk storages and rendering speed are presented in Table IX. It can be seen that there exists trade-offs between the rendering performance and the texture resolution. By choosing 2048 as the resolution, we obtain less disk storages without compromising rendering quality and rendering speed too much.
TABLE IX: Ablation of texture resolution. We report PSNR, disk storages and rendering speed for reference
Resolution PSNR $\uparrow$ Disk Storage $\downarrow$ Rendering Speed $\uparrow$
2048 27.58 16.5 MB 50 FPS
1024 27.06 14.4 MB 54 FPS
4096 27.74 19.8 MB 44 FPS